OpenStudy (anonymous):

need help with fractions 5/6 + 1/12= please break down

OpenStudy (aroub):

5/6*2 +1/12 10/12+1/12 11/12

OpenStudy (lalaly):

\[\frac{5}{6} \times \frac{2}{2} + \frac{1}{12}= \]\[\frac{10}{12} + \frac{1}{12} = \frac{11}{12}\]

OpenStudy (anonymous):

trying to help my son with math home work. 9/14 + 2/7=

OpenStudy (anonymous):

trying to help my son with math home work. 9/14 + 2/7=

OpenStudy (lalaly):

same way u need a common denominator, so just multiply the 2/7 by 2/2 to make the denominator 14 then add the fractions